Ticket Number: HEA-NOC/20060207-2 Ticket Status: OPEN
Ticket Type: unscheduled Resolver: TCD
Ticket Opened: 20060207 17:18 UTC Problem Start: 20060207 16:15 UTC
Ticket Update: 20060207 17:18 UTC Problem End:
Internal routing problem in TCD.
Traffic to and from TCD is effected.
20060207 17:03 OMG Rang Brian O'Hora to notify him that other clients were
having trouble reaching TCD. There was a drop off in traffic on the MRTG
graphs at this time. Troubleshooting seems to indicate that they have lost
a default route from gw-18.
This ticket can be monitored at http://www.hea.net/tickets/20060207-2